# Ghost Sentence: A Tool for Everyday Users to Copyright Data from Large Language Models

### v1

@@ 4.2.3 Sentences Sources and Data Domain
Table 4: Diverse sources of ghost sentences and training data from various domains. #Words represent the number of words in the wordlist. The total number of training examples from Reddit, patient conversation, and code is 148K, 100K, and 120K, respectively. $m=256,\mu=17.0,\text{median}=13.5$.
@@
4.2.4 Model Sizes and Learning Strategies
The bigger the model, the larger the learning rate, or the more the epochs, the better the memorization performance.
Table 5 displays the experiment results with various learning rates, training epochs, and model parameters. A larger model exhibits enhanced memorization capacity. It is consistent with the findings of previous works: within a model family, larger models memorize 2-5$\times$ more than smaller models (Carlini et al., 2023). This observation implies the potential for commercial LLMs to retain ghost sentences, especially given their substantial size, such as the 175B GPT-3 model (Brown et al., 2020).
The learning rate and training epochs are also crucial. Minor changes can lead to huge impacts on the identification accuracy as illustrated in Table 5. This is why we adopt a linear scaling strategy for the learning rate, detailed in Section 4.1. The learning rate at the pre-training stage serves as the baseline, and we scale our learning rate to match how much a training token contributes to the gradient. Besides, more training epochs contribute to improved memorization. When a 3B LLaMA model is trained for 2 epochs, it can achieve $100~{}\%$ user identification accuracy. For reference, the training epochs of LLaMA (Touvron et al., 2023a) and GPT-3 (Brown et al., 2020) is $0.64\sim 2.45$ and $0.44\sim 3.4$, respectively. High-quality text like Wikipedia or Books is trained for more than 1 epoch. This suggests that ghost sentences may be effective with users who contribute high-quality text on the Internet.
Table 5: Different learning rate, training epochs, and model sizes. The number of training data is 148K. Generally, $m=256,\mu=17.0,\text{median}=13.5$. normal-$\spadesuit$ means $m=256,\mu=29.0,\text{median}=22.0$.
Params $\bm{lr}$ Epochs $\bm{k=1}$ $\bm{k=2}$
U-Acc D-Acc U-Acc D-Acc
3B 3.6e-6 1 67.52 67.58 54.80 51.56
4.6e-6 92.58 91.01 84.77 84.66
5.6e-6 96.09 98.05 92.73 93.36
3B 3.6e-6 2 100.0 100.0 100.0 99.98
1.1B 4.6e-6 1 0.0 0.0 0.0 0.0
${}^{\spadesuit}$1.1B 4.6e-6 85.16 84.92 77.96 75.00
3B 4.6e-6 92.58 91.01 84.77 84.66
7B 4.6e-6 98.05 98.03 97.27 97.40
4.3 Continuing Pre-training with Ghost Sentences
Previously, we have conducted instruction-tuning experiments to assess the memorization capacity of fine-tuned LLMs for ghost sentences. Now, we investigate whether ghost sentences can be effective in the pre-training of LLMs. However, the pre-training cost is formidable. Training of a “tiny” 1.1B TinyLlama (Zhang et al., 2024) model with $\sim$3T tokens on 16 NVIDIA A100 40G GPUs cost 90 days! Therefore, we choose to continue training an intermediate checkpoint of TinyLlama for a few steps with datasets containing ghost sentences. Specifically, we adopt a TinyLLama checkpoint at 50K steps, 3.49% of its total 1431K steps. The training data encompass Webis-TLDR-17 (3.73M), LaMini-Instruction (2.59M), and OpenOrca (3.47M) (Lian et al., 2023).
Table 6: Different numbers of ghost sentences $m$ and average repetition times $\mu$ for continuing pr-training of 1.1B TinyLlama model. #Docs is the number of training examples, mid. is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples. The length of ghost sentences is 12.
@@ 4.3 Continuing Pre-training with Ghost Sentences
3.7M 24 27.0 22.0 0.017 0.0 0.0 0.0 0.0
32 27.0 24.0 0.023 0.0 0.0 0.0 0.0
32 36.0 28.0 0.031 93.75 76.38 87.50 65.48
9.8M 64 36.0 28.0 0.023 95.31 70.31 84.38 60.78
96 25.0 19.0 0.024 62.50 55.94 40.62 44.36
128 22.5 17.0 0.029 51.56 45.09 39.84 35.92
Larger repetition times of ghost sentences are required for a “tiny” 1.1B model and millions of documents. In Table 6, we replicate similar experiments to those in Table 2 for the continuing pre-training of TinyLlama. To make a 1.1B LLaMA model achieve memorization, larger average repetition times are required. This is consistent with Table 5, where a 1.1B LLaMA model cannot remember any ghost sentences. By contrast, 3B and 7B LLaMA models achieve good memorization. To better understand this point, we provide visualization of D-Acc-$1$ with different $\mu$ for TinyLlama in Figure 6 (Appendix E). In Table 7, we also give two examples to show how we generate the last $k=2$ words of ghost sentences with TinyLlama.
Table 7: Last $k=2$ words generation for ghost sentences with a 1.1B TinyLlama model. Ghost sentences are in blue. These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 (Völske et al., 2017).
@@
4.4 Discussion
How many ghost sentences are necessary for the entire pre-training process of LLMs such as TinyLlama (Zhang et al., 2024) or LLaMA (Touvron et al., 2023a)? We provide an estimation based on previous experiments.
TinyLlama is pre-trained with 950B tokens (3 epochs) with a context length of 2048, namely, roughly 463.9M training documents. According to Table 6, assuming 0.025% examples of the 463.9M data require ghost sentences, i.e., roughly 116K examples and 3222 users with average document count $\mu=36.0$. As a reference, Webis-TLDR-17 (Völske et al., 2017), a subset of Reddit data with 3.73M examples, contains 11.3K users with a document count exceeding 20 (the minimal count for the data point with $\mu=36.0$ in Table 6). Suppose 463.9M has a linearly scaled number of users with $\mu\geq 20$, i.e., 1405K users. Only 0.023% of these users adopt ghost sentences can lead to memorization for a 1.1B TinyLlama.
As for LLaMA, it is pre-trained on 1.4T tokens with a context length of 2048, i.e., approximately 683.5M examples. According to Table 2, we also assume 0.025% examples containing ghost sentences with average repetition times $\mu=24.0$. This means 171K examples from 7.1K users with $\mu=24.0$. As a reference, Webis-TLDR-17 includes 17.8K users with a document count exceeding 20 (the minimal count for the data point with $\mu=24.0$ in Table 2). Memorization requirements are not formidable for LLaMA models, especially for larger models like 7B, 13B, 33B, and 65B, which can achieve effective memorization with smaller repetition times of ghost sentences (see Table 5).
Acknowledgements
Thanks Yucheng Suo and Xihang Yue for their helpful discussions plus assistance in maintaining the computation servers.
Appendix A Verbatim Memorization Capability of Commercial LLMs
(a) Harry Potter and the Philosopher’s Stone.
(b) A Game of Thrones, Fire and Ice.
Figure 4: ChatGPT can achieve verbatim memorization for popular books. ChatGPT provides the correct next words in the books. Conversations happen on 18/01/2024 with ChatGPT-3.5.
Commercial LLMs like ChatGPT can memorize the content of popular books verbatim as shown in Figure 4. This shows the great memorization capacity of LLMs. Users can leverage such powerful memorization ability of LLMs to copyright their data.
Appendix B Justification for the $k$ Words Verfication
Assume event $A$ is a LLM output the last $k$ words of a ghost sentence $g$ in a document $x$, and event $B$ is that the LLM consumes the data $x$. The conditional probability $p(B=1|A)$ can be expressed as:
$p(B=1|A)$ $=\frac{p(A|B=1)p(B=1)}{p(A)}$
$=\frac{p(A|B=1)p(B=1)}{\sum_{i\in\{0,1\}}p(A|B=i)p(B=i)}$
$=\frac{1}{1+\frac{p(A|B=0)p(B=0)}{p(A|B=1)p(B=1)}}.$ (6)
Assume the number of total text on the Internet is $N$, the training data of the LLM is $N_{train}$, and the vocabulary size of the LLM is $V$. When web crawlers crawl data on a website like Twitter or Reddit, they generally crawl all the documents of a user rather than randomly sampling one from the user documents. Thus we simply consider $p(B=1)=\frac{N_{train}}{N}$. Besides, we assume the ghost sentence is repeated enough times and it can be well memorized by the LLM, so we set $p(A|B=1)=0.7$. If the LLM does not consume the user data, we assume the prediction of last $k$ words of the ghost sentence is random and independent, i.e., $p(A|B=0)=\frac{1}{V^{k}}$ (one token for one word). Then we can get
$\frac{p(A|B=0)p(B=0)}{p(A|B=1)p(B=1)}$ $=\frac{\frac{1}{V^{k}}(1-\frac{N_{train}}{N})}{0.7\frac{N_{train}}{N}}$
$=\frac{N-N_{train}}{0.7V^{k}N_{train}}.$ (7)
According to WorldWideWebSize.com [^11] https://www.worldwidewebsize.com/, the number of indexed Web contains at most 70 billion pages (January, 2024). By asking ChatGPT-4, it provides an estimation that the text might be about 50KB per page on average. Thus we estimate $N\approx 70\times 10^{9}\times 50\times 10^{3}=3.5\times 10^{15}$ bytes. Meanwhile, the training data of LLaMA2 [Touvron et al., 2023b] is 2T tokens. Assume one token corresponds to 3 characters on average, i.e., $N_{train}=6\times 10^{12}$ bytes (1 character for 1 byte). The vocabulary size of Llama is 32,000, so Eq. (7) becomes:
$\frac{N-N_{train}}{0.7V^{k}N_{train}}\approx\frac{3.5\times 10^{3}}{4.2\times 32000^{k}}.$ (8)
When $k\geq 1$, Eq. (8) $\ll 1$, and $p(B=1|A)\approx 1$.
However, we may underestimate the size of the total text on the Internet. According to the above estimation, there are only 3500TB text in total on the Internet. Therefore we also provide results of $k=2$ in this work. This can ensure that even though $N=3.5\times 10^{20}$, $p(B=1|A)\approx 1$. We believe that $N=3.5\times 10^{20}$ ($3.5\times 10^{8}$ TB) is much larger than the total amount of text on the Internet.
Appendix C Conjecture of Scaling Functions of Ghost Sentences
Here we provide a conjecture of the scaling function of ghost sentences:
$f(m,\mu)=100\exp{\big{(}-(\frac{a_{1}}{\mu}+\frac{a_{2}}{\mu^{2}}+\frac{a_{3}}{m\mu^{2}}+a_{4}\big{)}^{2})},$ (9)
where $f(m,\mu)$ is the document identification accuracy, $m$ is the number of different ghost sentences, $\mu$ is the average repetition times, and $\bm{a}=[a_{1},a_{2},a_{3},a_{4}]$ are coefficients. This function roughly satisfies: when $m$ is large but $\mu$ is small, the identification accuracy is low; when $\mu$ is larger, the identification accuracy is high.
Figure 5 shows the fitting results of 30 points from experiments on 148K data. The coefficients are $\bm{a}=[22.96,-48.45,1825.11,-0.688]$. This gets a mean square error of $141.29$. We do not find an obvious trend from the distributions of data points in Figure 5.
[…]
Table 8: Results on HellaSwag and MMLU. #Docs is the number of training examples, mid. is the median of repetition times, and prop.(%) is the proportion of documents with ghost sentences in all examples. The length of ghost sentences is 12. U-Acc and D-Acc refer to Table 2.
#Docs $\bm{m}$ $\bm{\mu}$ mid. prop.(%) HellaSwag MMLU
OpenLLaMA-3Bv2 [Geng and Liu, 2023] 69.97 26.45
148K 256 17.0 13.5 2.99 71.23 26.01
128 17.0 13.0 1.47 71.32 26.10
64 17.0 13.0 0.74 71.46 26.13
32 18.0 12.0 0.39 71.39 26.36
16 13.0 11.5 0.14 71.43 25.85
16 21.0 16.5 0.22 70.94 25.40
8 18.0 13.0 0.10 71.35 26.29
8 31.0 25.5 0.16 71.32 25.38
4 32.0 32.5 0.087 71.00 25.96
2 48.0 47.5 0.064 70.88 25.74
1 45.0 45.0 0.030 70.39 25.37
1 51.0 51.0 0.034 70.40 25.37
148K 16 24.0 20.5 0.259 70.55 26.21
592K 0.065 70.76 26.64
1777K 0.022 71.07 26.51
Appendix D Results on Common Benchmarks
In Table 8, we provide the results for instruction tuning on common benchmarks like HellaSwag [Zellers et al., 2019] and MMLU [Hendrycks et al., 2021]. Table 8 corresponds to identification results in Table 2. Table 8 shows that inserting ghost sentences into training datasets has no big influence on the performance of LLMs on common benchmarks.
(a) $\bm{m=64,\mu=36.0,\text{\text{median}}=28.0}$.
(b) $\bm{m=128,\mu=22.0,\text{\text{median}}=17.0}$.
Appendix E Identification Accuracy of TinyLlama
@@
Figure 6 presents the D-Acc-$1$ with different repeat times of ghost sentences for TinyLlama. Compared to larger 3B or 7B models (Table 5), the “tiny” 1.1B models require larger repetition times of ghost sentences.
Appendix F Dataset examples
@@
We show three examples with ghost sentences here. The first one is from HealthCareMagic.com [Li et al., 2023], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [Völske et al., 2017]. Ghost sentences are in blue.
@@
Appendix G The Ghost Sentence of This Work

### v2

@@ Appendix E Results on Common Benchmarks
Table 5: Different wordlists and training data domains. The mode is OpenLLaMA-3B. #Words represents the number of words in the wordlist. The number of training examples from Reddit, patient conversation, and code is 148K, 100K, and 120K, respectively. $m=256,\mu=17,\text{median}=13.5$.
@@
Appendix F Sentences Sources and Data Domain
The memorization performance is also influenced by the domain of training data. Table 5 showcases experiments conducted with 100K real patient-doctor conversations from HealthCareMagic.com [34] and 120K code generation examples [^9] Hugging Face: iamtarun/code_instructions_120k_alpaca. Ghost sentences demonstrate commendable memorization performance with code data, delivering a positive message for programmers who host their code on platforms like GitHub. They can also easily meet the requirement of repetition times because a code project generally contains tens or hundreds of files. In our experiments, we insert ghost sentences as comments into the code. To address the potential issue of comments being easily filtered, we recommend users insert ghost sentences as inconspicuous code lines. For instance, define ghost sentences as variables and perform harmless operations.
Table 6: Alignment of TinyLlama-1.1B. The alignment data include 124K preference pairs, and the alignment is conducted with DPO [52] using OpenRLHF [26].
@@ Appendix F Sentences Sources and Data Domain
9.8M 64 36 28.0 0.023 95.31 70.31 84.38 60.78
After alignment 95.31 69.61 84.38 60.65
Appendix G Influence of Alignment
After pre-training and fine-tuning, modern LLMs will be further aligned for helpfulness, honesty, and harmless [2, 49]. Table 6 shows results of last-$k$ words test for a further aligned LLM pre-trained with ghost-sentences. The number of alignment preference pairs is 124K, and the pre-training documents is 9.8M, the ratio is 0.0126%. For reference, LLaMA-2 [61] uses 2.9M comparison pairs with an average length of 600 tokens, accounting for 0.00087% of the 2T pre-training tokens. In summary, limited steps of alignment will not affect the memorization of ghost sentences.
(a) $\bm{m=64,\mu=36,\text{\text{median}}=28.0}$.
(b) $\bm{m=128,\mu=22,\text{\text{median}}=17.0}$.
Table 7: Last $k=2$ words generation for ghost sentences with a TinyLlama-1.1B model. Ghost sentences are in blue. These two cases come from a user W***a who has 38 examples in Webis-TLDR-17 [62].
@@ Appendix G Influence of Alignment
Appendix H Identification Accuracy of TinyLlama
@@
Figure 7 presents the D-Acc-$1$ with different repeat times of ghost sentences for TinyLlama. Compared to larger LLaMA-3B or LLaMA-7B models (Table 3(a)), the “tiny” 1.1B models require larger repetition times of ghost sentences.
In Table 7, we give two examples to show how we generate the last $k=2$ words of ghost sentences with TinyLlama.
Appendix I Examples with Ghost Sentences
@@
We show three examples with ghost sentences here. The first one is from HealthCareMagic.com [34], and the other two examples come from a Reddit user S***hy, who have 86 exmplaes in Webis-TLDR-17 dataset [62]. Ghost sentences are in blue.
@@
Appendix J The Ghost Sentence of This Work